OpenStudy (anonymous):

4. A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 60t + 6. What will the height be at t = 3 seconds?

OpenStudy (anonymous):

replace t by 3 and (carefully) compute

OpenStudy (anonymous):

\[h(t) = –16t² + 60t + 6\] \[h(3) = –16\times 3^2 + 60\times 3 + 6\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so whats the answer